OpenStudy (anonymous):

You decide to put $150 in a savings account to save for a $3,000 down payment on a new car. If the account has an interest rate of 2.5% per year and is compounded monthly, how long does it take you to earn $3,000 without depositing any additional funds?

OpenStudy (anonymous):

Use the compound interest formula \[F=P \left( 1+\frac{ r }{ n } \right)^{nt}\] F = future amount, $3000 P = present amount, $150 r = interest rate in decimal, 0.025 n = number of compoundings per year, 12 t = number of years

OpenStudy (anonymous):

@kaykay23

OpenStudy (anonymous):

how would I get the t in this situation, once I have gotten 3000=150(1.002)^12t

OpenStudy (anonymous):

first divide both sides by 150, then convert to a log equation.

OpenStudy (anonymous):

\[20=(1.0020833)^{12t}\] \[\log 20 = \log (1.0020833)^{12t}\] \[\log 20=12t \log1.0020833\]

OpenStudy (anonymous):

how would I go on?